you drop a stone in a well and hear it hit the bottom in 6.19s how deep is the well
1
Expert's answer
2019-03-19T12:06:14-0400

Let the depth of well is L. The total time

"t=t_1+t_2=\sqrt{\frac{2L}{g}}+\frac{L}{c}"

So, we have an equation

"6.19=\sqrt{\frac{2L}{9.8}}+\frac{L}{340}"

Solution


"L=160\:\rm{m}"